# Existence and uniqueness of solutions to SDEs with state-dependent variable exponent

### v1

@@ Abstract
We study a time-inhomogeneous, nonlinear stochastic differential equation whose drift and diffusion terms involve state-dependent variable exponents. This novel structure introduces flexibility, allowing the system to adapt dynamically to its current state and better capture real-world phenomena compared to classical models such as GBM and CEV. While this flexibility enhances modeling accuracy, it also creates analytical challenges. Using a fixed-point approach, we establish existence and uniqueness under suitable conditions. The model generalizes GBM and CEV while exhibiting distinct behavior for small states.
@@
Stochastic process; state-dependent variable exponent; Banach fixed-point theorem.
@@ 1 Introduction
where $W(t)$ is a Brownian motion on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$; $p(\cdot),q(\cdot)$ are differentiable, variable exponent functions; and $0<\mu,\sigma$ are bounded real-valued deterministic functions.
@@
In this paper, building on the research findings of [1, 2, 3], we differently study a time-inhomogeneous model, and obtain the existence and uniqueness of solution by exploiting the idea of the contraction mapping principle, also known as the Banach fixed-point theorem.
@@ 2 Preliminaries and auxiliary results
    $\mu,\sigma:[0,T]\to(0,\infty)$ are deterministic functions such that
    $0<\mu^{-}:=\inf_{t\in[0,T]}\mu(t),\quad\sup_{t\in[0,T]}\mu(t):=\mu^{+}<\infty.$
    $0<\sigma^{-}:=\inf_{t\in[0,T]}\sigma(t),\quad\sup_{t\in[0,T]}\sigma(t):=\sigma^{+}<\infty.$
@@
Let $\{\mathcal{F}_{t}\}_{t\geq 0}$ be a filtration denoted by $\mathcal{F}_{t}$. Then $(\Omega,\mathcal{F},\mathcal{F}_{t},\mathbb{P})$ is a filtered probability space satisfying the usual conditions (see, e.g., [8, 9]). A stochastic process $X(t)$, defined for $t\geq 0$, is said to be progressively measurable with respect to a filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ if, for every $t\in[0,T]$, the mapping
@@
It is enough to apply ($\mathbf{h_{2}}$). If $0<x\leq 1$, there exists $0<\varepsilon\leq 1$ such that
$x^{h(x)}\leq\varepsilon^{h^{-}}\leq\varepsilon^{h^{-}}(1+x).$ (4)
@@
where $K:=\max\{\varepsilon^{h^{-}},e^{M_{\infty}}\}$. Hence, considering all these along with the fact that $X(t)\in\mathcal{M_{T}}$, it follows that
$\mathbb{E}\left[\int_{0}^{T}|X(t)|^{2h(X(t))}dt\right]\leq 2K^{2}(T+1)\mathbb{E}\left[\int_{0}^{T}X^{2}(t)ds\right]=2K^{2}(T+1)\|X\|^{2}_{\mathcal{T}}<\infty.$ (8)
Lemma 2.2.
@@ 3 The main result
Assume that $p(\cdot),q(\cdot)\in\mathcal{S}$ and condition $(\mathbf{D})$ holds. Also, suppose the initial condition $X_{0}=X(0)=x_{0}$ is $\mathcal{F}_{0}$-measurable, independent of $W(t)$, and has a finite second moment. Then there exists a unique solution $X(t)$ of ($\mathcal{GM}$}) with continuous paths satisfying $\mathbb{E}\left[\int_{0}^{T}|X(t)|^{2}dt\right]<\infty$.
@@
$\leq 3\mathbb{E}[x_{0}^{2}]+6\mu^{2}_{+}K^{2}(T+1)\mathbb{E}\left[\int_{0}^{t\wedge\tau_{n}}X^{2}_{n}(s)ds\right]+3\sigma^{2}_{+}\mathbb{E}\left[\int_{0}^{t}\mathbf{1}_{[0,\tau_{n}]}(s)X^{2q}_{n}(s)ds\right]$
@@
where $c_{1}=3\mathbb{E}[x_{0}^{2}]$ and $c_{2}=3[2(T+1)K^{2}(\mu^{2}_{+}+\sigma^{2}_{+})]$ are independent of $n$. If we define $f_{n}(t)=\mathbb{E}[X^{2}_{n}(t)]$, and apply the Gronwall lemma to (3) we obtain
$f_{n}(T)\leq c_{1}e^{c_{2}T}=c_{3}\Longrightarrow\sup_{n}f_{n}(T)<\infty\,\,\text{as }n\to\infty.$ (26)

### v2

# On the existence, uniqueness and stability of solutions of SDEs with state-dependent variable exponent

@@ Abstract
We study a time-inhomogeneous nonlinear SDE with drift and diffusion governed by state-dependent variable exponents. This framework generalizes models like the geometric Brownian motion (GBM) and the constant elasticity of variance (CEV), offering flexibility to capture complex dynamics while posing analytical challenges. Using a fixed-point approach, we prove existence and uniqueness, analyze higher-order moments, derive asymptotic estimates, and assess stability. Finally, we illustrate an application where Poisson’s equation admits a probabilistic representation via a time-homogeneous nonlinear SDE with state-dependent variable exponents.
@@
Stochastic process; state-dependent variable exponent; higher-order moments; asymptotic estimate; stability; Banach fixed-point theorem; Feynman–Kac formula.
@@ 1 Introduction
where $W(t)$ is a Brownian motion on the complete probability space $(\Omega,\mathcal{F},\mathbb{P})$; $p(\cdot),q(\cdot)$ are differentiable, variable exponent functions; and $0<\mu,\sigma$ are bounded real-valued deterministic continuous functions.
@@
In this paper, building on the research findings of [1, 2, 3], we differently study a time-inhomogeneous model, and obtain the existence and uniqueness of solution by exploiting the idea of the contraction mapping principle, also known as the Banach fixed-point theorem. Additionally, we investigate higher-order moments ($\geq$ second), asymptotic estimates, and the stability of the solution. Lastly, we provide an application where a probabilistic representation of the solution of Poisson’s equation is given via the solution of a time-homogeneous nonlinear stochastic differential equation with state-dependent variable exponent.
@@ 2 Preliminaries and auxiliary results
    $\mu,\sigma:[0,T]\to(0,\infty)$ are deterministic continuous functions such that
    $0<\mu^{-}:=\min_{t\in[0,T]}\mu(t),\quad\max_{t\in[0,T]}\mu(t):=\mu^{+}<\infty.$
    $0<\sigma^{-}:=\min_{t\in[0,T]}\sigma(t),\quad\max_{t\in[0,T]}\sigma(t):=\sigma^{+}<\infty.$
@@
Let $\{\mathcal{F}_{t}\}_{t\geq 0}$ be a filtration denoted by $\mathcal{F}_{t}$. Then $(\Omega,\mathcal{F},\mathcal{F}_{t},\mathbb{P})$ is a filtered probability space satisfying the usual conditions (see, e.g., [8, 9]). A stochastic process $X(t):=\{X(t)\}_{0\leq t\leq T}$, defined for $t\geq 0$, is said to be progressively measurable with respect to a filtration $\mathcal{F}_{t}$ if, for every $t\in[0,T]$, the mapping
@@
It is enough to apply ($\mathbf{h_{2}}$). If $0<x\leq 1$, there exists $0<\epsilon\leq 1$ such that
$x^{h(x)}\leq\epsilon^{h^{-}}\leq\epsilon^{h^{-}}(1+x).$ (4)
@@
where $K:=\max\{\epsilon^{h^{-}},e^{M_{\infty}}\}$. Hence, considering all these along with the fact that $X(t)\in\mathcal{M_{T}}$, it follows that
$\mathbb{E}\left[\int_{0}^{T}X^{2h}(t)dt\right]\leq 2K^{2}(T+1)\mathbb{E}\left[\int_{0}^{T}X^{2}(t)ds\right]=2K^{2}(T+1)\|X\|^{2}_{\mathcal{T}}<\infty.$ (8)
Lemma $2.$
@@
Lemma 2.3.
[8] Let $m\geq 2$ be a real number. Let $f\in\mathcal{M_{T}}$ such that
$\mathbb{E}\left[\int_{0}^{T}|f(s)|^{m}ds\right]<\infty.$ (12)
Then
$\mathbb{E}\left[\left|\int_{0}^{T}f(s)dW(s)\right|^{m}\right]\leq\left(\frac{m(m-1)}{2}\right)^{\frac{m}{2}}\,T^{\frac{m-2}{2}}\,\mathbb{E}\left[\int_{0}^{T}|f(s)|^{m}ds\right].$ (13)
In particular, for $m=2$, there is equality, i.e. (13) gives Itô isometry.
Lemma 2.4.
[8] Let $\alpha,\beta,T$ be any positive numbers. Let $g\in\mathcal{L}_{2}([0,T];\mathbb{R})$ (i.e. the family of $\mathbb{R}$-valued $\mathcal{F}_{t}$-adapted processes) such that
$\mathbb{P}\left\{\int_{t_{0}}^{T}|g(t)|^{2}dt<\infty\right\}=1\quad\text{for all}\quad 0\leq t_{0}<T.$ (14)
Then
$\mathbb{P}\left\{\sup_{t_{0}\leq t\leq T}\left[\int_{t_{0}}^{t}g(s)dW(s)-\frac{\alpha}{2}\int_{t_{0}}^{t}|g(s)|^{2}ds\right]>\beta\right\}\leq e^{-\alpha\beta}.$ (15)
@@ 3 The main result
Assume that $p(\cdot),q(\cdot)\in\mathcal{S}$ and condition $(\mathbf{D})$ holds. Also, suppose the initial condition $X_{0}=X(0)=x_{0}$ is $\mathcal{F}_{0}$-measurable, independent of $W(t)$, and has a finite second moment. Then there exists a unique global solution $X(t)$ of ($\mathcal{GM}$}) with continuous paths satisfying $\mathbb{E}\left[\int_{0}^{T}|X(t)|^{2}dt\right]<\infty$.
@@
where $c_{1}=3\mathbb{E}[x_{0}^{2}]$ and $c_{2}=3[2(T+1)K^{2}(\mu^{2}_{+}+\sigma^{2}_{+})]$ are independent of $n$. If we define $\varphi_{n}(t):=\mathbb{E}[X^{2}_{n}(t)]$, and apply the Gronwall lemma to (3) we obtain
$\varphi_{n}(T)\leq c_{1}e^{c_{2}T}=c_{3}\Longrightarrow\sup_{n}\varphi_{n}(T)<\infty\,\,\text{as }n\to\infty.$ (30)
@@
4 The properties of the solution
In this section, by assuming that $X(t)$, with the initial value $X(t_{0})=x_{0}$, $0\leq t_{0}\leq T$, where $x_{0}$ is a $\mathcal{F}_{t_{0}}$-measurable random variable, is the unique solution for the equation ($\mathcal{GM}$}), we establish some important properties of the solution.
4.1 The higher-order moments of the solution
Theorem 4.1.
Let $m\geq 2$ be a real number and $\mathbb{E}\left[x_{0}^{m}\right]<\infty$. Then
$\mathbb{E}\left[X^{m}(t)\right]$ $\leq\left(3^{m-1}\mathbb{E}[x_{0}^{2}]+(t-t_{0})(A+B)\right)\,e^{t(A+B)},$ (32)
where $A=6^{m-1}(\mu^{+})^{m}(t-t_{0})^{m-1}K^{m}$ and $B=6^{m-1}(\sigma^{+})^{m}\left(\frac{m(m-1)}{2}\right)^{\frac{m}{2}}\,(t-t_{0})^{\frac{m-2}{2}}\,K^{m}$.
Proof.
Let’s first define the stopping time $\hat{\tau}_{n}$ by
$\hat{\tau}_{n}=T\wedge\inf\{t\geq 0;\,X_{n}(t)\notin(0,n),\,\,n\in\mathbb{N}\},$ (33)
from which we have $\hat{\tau}_{n}\uparrow T$ $\mathbb{P}$-a.s. Set $X_{n}(t)=X(t\wedge\hat{\tau}_{n})$ for $t\in[0,T]$. Then proceeding as before but this time using Lemma 2.3 and the linear growth (7) provides
$\mathbb{E}\left[X^{m}_{n}(t)\right]$ $\leq 3^{m-1}\mathbb{E}[x_{0}^{2}]+3^{m-1}(\mu^{+})^{m}\mathbb{E}\left[\left(\int_{t_{0}}^{t\wedge\hat{\tau}_{n}}X_{n}^{p}(s)ds\right)^{m}\right]$
$+3^{m-1}(\sigma^{+})^{m}\mathbb{E}\left[\left(\int_{t_{0}}^{t\wedge\hat{\tau}_{n}}X_{n}^{q}(s)dW(s)\right)^{m}\right]$
$\leq 3^{m-1}\mathbb{E}[x_{0}^{2}]+3^{m-1}(\mu^{+})^{m}(t-t_{0})^{m-1}\mathbb{E}\left[\int_{t_{0}}^{t\wedge\hat{\tau}_{n}}X^{mp}_{n}(s)ds\right]$
$+3^{m-1}(\sigma^{+})^{m}\left(\frac{m(m-1)}{2}\right)^{\frac{m}{2}}\,(t-t_{0})^{\frac{m-2}{2}}\,\mathbb{E}\left[\int_{t_{0}}^{t\wedge\hat{\tau}_{n}}X^{qm}_{n}(s)ds\right]$
$\leq 3^{m-1}\mathbb{E}[x_{0}^{2}]+(t-t_{0})(A+B)+(A+B)\int_{t_{0}}^{t\wedge\hat{\tau}_{n}}\mathbb{E}\left[X^{m}_{n}(s)\right]ds$
$\leq 3^{m-1}\mathbb{E}[x_{0}^{2}]+(t-t_{0})(A+B)+(A+B)\int_{t_{0}}^{t}\mathbb{E}\left[X^{m}(s\wedge\hat{\tau}_{n})\right]ds.$ (34)
Applying the Gronwall inequality gives
$\mathbb{E}\left[X^{m}(t\wedge\hat{\tau}_{n})\right]$ $\leq\left(3^{m-1}\mathbb{E}[x_{0}^{2}]+(t-t_{0})(A+B)\right)\,e^{t(A+B)}.$ (35)
Finally, letting $n\to\infty$, and considering that $\hat{\tau}_{n}\to T$ as $n\to\infty$ $\mathbb{P}$-a.s., gives
$\mathbb{E}\left[X^{m}(t)\right]$ $\leq\left(3^{m-1}\mathbb{E}[x_{0}^{2}]+(t-t_{0})(A+B)\right)\,e^{t(A+B)}.$ (36)
4.2 The asymptotic estimate of the solution
Theorem 4.2.
Assume that the linear growth condition (7) holds. Then the solution $X(t)$ of ($\mathcal{GM}$}) $\mathbb{P}$-a.s. has the following property
$\limsup_{t\to\infty}\log(X(t))\leq\hat{K}$, (37)
where $\hat{K}=4\max\{K^{2}(\sigma^{+})^{2},K\mu^{+}\}$.
Proof.
Applying Itô’s formula gives
$d(\log(1+X^{2}(t)))$ $=\frac{1}{1+X^{2}(t)}\left[2X(t)\mu(t)X^{p}(t)+\sigma^{2}(t)X^{2q}(t)\right]dt$
$-\frac{2X^{2}(t)\sigma^{2}(t)X^{2q}(t)}{(1+X^{2}(t))^{2}}dt+\frac{2X(t)\sigma(t)X^{q}(t)}{1+X^{2}(t)}dW(t).$ (38)
Note that from the linear growth (7), one can obtain
$x\mu(t)x^{p(x)}+\frac{1}{2}\sigma^{2}(t)x^{2q(x)}\leq\hat{K}(1+x^{2}).$ (39)
Then
$\log(1+X^{2}(t))$ $=\log(1+x^{2}_{0})+2\hat{K}(t-t_{0})$
$-2\int_{t_{0}}^{t}\frac{(\sigma(s)X(s)X^{q}(s))^{2}}{(1+X^{2}(s))^{2}}ds+2\int_{t_{0}}^{t}\frac{\sigma(s)X(s)X^{q}(s)}{1+X^{2}(s)}dW(s).$ (40)
Then, for any integer $n\geq t_{0}$, using Lemma 2.4 gives
$\mathbb{P}\left\{\sup_{t_{0}\leq t\leq n}\left[2\int_{t_{0}}^{t}\frac{\sigma(s)X(s)X^{q}(s)}{1+X^{2}(s)}dW(s)-2\int_{t_{0}}^{t}\frac{(\sigma(s)X(s)X^{q}(s))^{2}}{(1+X^{2}(s))^{2}}ds\right]>2\log(n)\right\}\leq\frac{1}{n^{2}}.$ (41)
Define the events
$E_{n}=\left\{\sup_{t_{0}\leq t\leq n}\left[\int_{t_{0}}^{t}\frac{\sigma(s)X(s)X^{q}(s)}{1+X^{2}(s)}dW(s)-\int_{t_{0}}^{t}\frac{(\sigma(s)X(s)X^{q}(s))^{2}}{(1+X^{2}(s))^{2}}ds\right]>\log(n)\right\}.$ (42)
Then
$\sum_{n=1}^{\infty}\mathbb{P}(E_{n})\leq\sum_{n=1}^{\infty}\frac{1}{n^{2}}<\infty,$ (43)
which, by the Borel-Cantelli lemma, implies
$\mathbb{P}(\limsup_{n\to\infty}E_{n})=\mathbb{P}(E_{n}\,\text{i.o})=0.$ (44)
Therefore, for almost all $\omega\in\Omega$ there exists a random integer $n_{0}=n_{0}(\omega)\geq t_{0}+1$ such that
$\sup_{t_{0}\leq t\leq n}\left[\int_{t_{0}}^{t}\frac{\sigma(s)X(s)X^{q}(s)}{1+X^{2}(s)}dW(s)-\int_{t_{0}}^{t}\frac{(\sigma(s)X(s)X^{q}(s))^{2}}{(1+X^{2}(s))^{2}}ds\right]\leq\log(n),\quad\text{if }\,n\geq n_{0}.$ (45)
Then for all $t_{0}\leq t\leq n$, $n\geq n_{0}$ $\mathbb{P}-a.s.$ we have
$\int_{t_{0}}^{t}\frac{\sigma(s)X(s)X^{q}(s)}{1+X^{2}(s)}dW(s)\leq\int_{t_{0}}^{t}\frac{(\sigma(s)X(s)X^{q}(s))^{2}}{(1+X^{2}(s))^{2}}ds+\log(n).$ (46)
Thus using (46) in (4.2) gives
$\log(1+X^{2}(t))$ $=\log(1+x^{2}_{0})+2\hat{K}(t-t_{0})+2\log(n).$ (47)
Hence, for almost all $\omega\in\Omega$, when $n\geq n_{0}$ and $n-1\leq t\leq n$, we can write
$\frac{1}{t}\log(1+X^{2}(t))$ $=\frac{1}{n-1}\left(\log(1+x^{2}_{0})+2\hat{K}(n-t_{0})+2\log(n)\right),$ (48)
from which one can, $\mathbb{P}$-a.s., obtain
$\limsup_{n\to\infty}\frac{1}{t}\log(X(t))$ $\leq\limsup_{n\to\infty}\frac{1}{2t}\log(1+X^{2}(t))$
$\leq\limsup_{n\to\infty}\frac{1}{2(n-1)}\left(\log(1+x^{2}_{0})+2\hat{K}(n-t_{0})+2\log(n)\right)=\hat{K}$. (49)
This completes the proof. ∎
4.3 The stability of the solution
Definition 1.
A solution $X_{\xi}(t)$ of ($\mathcal{GM}$}) with initial value $X(t_{0})=\xi$, $0\leq t_{0}\leq T$, is said to be stable in the $m$th power if for all $\varepsilon>0$ there exists $\delta>0$ such that when
$\mathbb{E}\left[|\xi-\eta|^{m}\right]<\delta$, then
$\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{\xi}(s)-X_{\eta}(s)|^{m}\right]$ $\leq\varepsilon,$ (50)
where $X_{\eta}(t)$ is another solution of ($\mathcal{GM}$}) with initial value $X(t_{0})=\eta$, and $\xi,\eta$ are $\mathcal{F}_{t_{0}}$-measurable random variables.
Theorem 4.3.
Let $X_{\xi}(t)$ and $X_{\eta}(t)$ be two solutions of ($\mathcal{GM}$}) starting from positions $\xi$ and $\eta$, respectively, where $\xi$ and $\eta$ are $\mathcal{F}_{t_{0}}$-measurable random variables. Then the solution of ($\mathcal{GM}$}) is stable in the $m$th power.
Proof.
We fix $t\in[0,T]$, and consider the sequence $x_{n}\to x$ $\mathbb{P}$-a.s. as $n\to\infty$ with $x\in(0,\infty)$, and $(x_{n})\subset(0,\infty)$. Let $X_{x_{n}}(t)$ and $X_{x}(t)$ denote the solutions of ($\mathcal{GM}$}) starting from positions $x_{n}$ and $x$, respectively. Then for any integer $m\geq 2$, it reads
$\sup_{0\leq s\leq t}|X_{x_{n}}(s)-X_{x}(s)|^{m}$ $\leq 3^{m-1}|x_{n}-x|^{m}$
$+3^{m-1}(\mu^{+})^{m}\sup_{0\leq s\leq t}\left|\int_{0}^{s}\left(X_{x_{n}}(r)^{p(X_{x_{n}}(r))}-X_{x}(r)^{p(X_{x}(r))}\right)dr\right|^{m}$
$+3^{m-1}(\sigma^{+})^{m}\sup_{0\leq s\leq t}\left|\int_{0}^{s}\left(X_{x_{n}}(r)^{q(X_{x_{n}}(r))}-X_{x}(r)^{q(X_{x}(r))}\right)dW(r)\right|^{m}.$ (51)
Applying the Burkholder-Davis-Gundy (BDG) inequality and the Hölder inequality together yields
$\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{x_{n}}(s)-X_{x}(s)|^{m}\right]$ $\leq 3^{m-1}\mathbb{E}\left[|x_{n}-x|^{m}\right]$
$+3^{m-1}(\mu^{+})^{m}L^{m}t^{m-1}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq s\leq r}\left|X_{x_{n}}(s)-X_{x}(s)\right|^{m}\right]dr$
$+3^{m-1}C_{m}\sigma^{+})^{m}L^{m}t^{(m-2)/2}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq s\leq r}\left|X_{x_{n}}(s)-X_{x}(s)\right|^{m}\right]dr$
$\leq 3^{m-1}\mathbb{E}\left[|x_{n}-x|^{m}\right]+\bar{L}\int_{0}^{t}\mathbb{E}\left[\sup_{0\leq s\leq r}\left|X_{x_{n}}(s)-X_{x}(s)\right|^{m}\right]dr.$ (52)
where $\bar{L}=3^{m-1}L^{m}((\mu^{+})^{m}t^{m-1}+C_{m}(\sigma^{+})^{m}t^{(m-2)/2})$. Lastly, applying the Gronwall inequality provides
$\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{x_{n}}(s)-X_{x}(s)|^{m}\right]$ $\leq 3^{m-1}\,e^{T\bar{L}}\,\mathbb{E}\left[|x_{n}-x|^{m}\right].$ (53)
Considering that $x_{n}\to x$ $\mathbb{P}$-a.s., there exists a number $\delta^{1/m}>0$ and a positive integer $n_{0}$ such that whenever $n\geq n_{0}$, it holds $|x_{n}-x|^{m}<\delta$, and hence, $\mathbb{E}\left[|x_{n}-x|^{m}\right]<\delta$. Thus, if we let $\delta=\frac{\varepsilon}{3^{m-1}\,e^{T\bar{L}}}$ in (53), it follows
$\mathbb{E}\left[\sup_{0\leq s\leq t}|X_{x_{n}}(s)-X_{x}(s)|^{m}\right]$ $\leq\varepsilon,$ (54)
which, by the Dominated Convergence theorem, implies
$\sup_{0\leq s\leq t}|X_{x_{n}}(s)-X_{x}(s)|\to 0\,\text{ in }\,L^{m}([0,T])\,\text{ whenever }\,x_{n}\to x\,\,\,\mathbb{P}-a.s.$ (55)
The proof is complete. ∎
5 Application
SDEs have numerous applications, one of the most notable being their role in representing solutions to partial differential equations (PDEs) through the Feynman–Kac formula. This formula establishes a fundamental connection between SDEs and PDEs, providing a probabilistic framework for analyzing and solving partial differential equations.
Let us consider Poisson’s equation
$\begin{cases}\begin{array}[]{rll}Lu(x)-cu(x)&=-f(x)\quad\text{in }\mathcal{O}=(a,b)\subset(0,\infty),\\ u(x)&=0\quad\text{on }\partial\mathcal{O},\end{array}\end{cases}$ (56)
where $f$ is a real-valued $C^{1}$ function; $0<\mu,\sigma$ and $c\geq 0$ real parameters; $L$ is the linear differential operator
$L=\frac{1}{2}\sigma^{2}x^{2q(x)}\frac{d^{2}}{dx^{2}}+\mu x^{p(x)}\frac{d}{dx}.$ (57)
Next, we show that for the coefficients of $L$, the following conditions are satisfied:
  • ($\mathbf{a_{1}}$)
    Since
    $\frac{1}{2}\sigma^{2}x^{2q(x)}\geq\frac{1}{2}\sigma^{2}\,\min\{a^{2q^{-}},a^{2q^{+}}\}:=\lambda,\,\,\lambda\in(0,\infty)$ (58)
    $L$ is uniformly elliptic in $\mathcal{O}$.
  • ($\mathbf{a_{2}}$)
    $\sigma^{2}x^{2q(x)}$ and $\mu x^{p(x)}$ are uniformly Lipschitz continuous on $\overline{\mathcal{O}}=[a,b]$.
    We’$l$ only provide a proof for $\vartheta(x):=x^{p(x)}$ since the same reasoning applies to $x^{2q(x)}$. Clearly $\vartheta$ is continuous on $\overline{\mathcal{O}}$. Further, using the assumptions ($\mathbf{h_{1}}$) and ($\mathbf{h_{3}}$), we obtain
    $|\vartheta^{\prime}(x)|\leq\max\{b^{p^{-}},b^{p^{+}}\}\left[\left(M_{0}+C_{0}a^{-(1+\alpha)}\right)\log(b)+\frac{p^{+}}{a}\right]$ (59)
    which implies, by the Mean Value theorem, that $\vartheta$ is uniformly Lipschitz continuous on $\overline{\mathcal{O}}$.
Under these assumptions, it is well-known that Poisson’s equation (56) admits a unique solution $u\in C^{2}(\mathcal{O})\cap C(\overline{\mathcal{O}})$.
Theorem 5.1.
Suppose $u$ is the unique solution to Poisson’s equation (56) in $\mathcal{O}$. Then $u$ is given by Feynman–Kac formula
$u(x)=\mathbb{E}^{x}\left[\int_{0}^{\tau_{\mathcal{O}}}e^{-cs}\,f(X(s))ds\right],$ (60)
where $\tau_{\mathcal{O}}=\inf\{t\geq 0:X(t)\notin\mathcal{O}\}$; $\mathbb{E}^{x}[\cdot]$ means expectation conditioned on $X_{0}=x$; and $X(t)$ is the unique global solution of time-homogeneous nonlinear SDE with state-dependent variable exponent
$\begin{cases}\begin{array}[]{rll}dX(t)&=\mu X^{p}(t)dt+\sigma X^{q}(t)dW(t),\,\,t\in[0,T],\\ X_{0}&=x\in\mathcal{O}.\end{array}\end{cases}$ (61)
Proof.
Since $u\in C^{2}$, we can apply Itô’s formula and obtain
$d(u(X(t))$ $=\left(\mu X^{p}(t)\frac{du(X(t))}{dx}+\frac{1}{2}\sigma^{2}X^{2q}(t)\frac{d^{2}u(X(t))}{dx^{2}}\right)dt$
$+\sigma X(t)^{q}\frac{du(X(t))}{dx}dW(t),$ (62)
and integrating over $[0,t]$
$u(X(t))$ $=u(X_{0})+\int_{0}^{t}\left(\mu X^{p}(s)\frac{du(X(s))}{dx}+\frac{1}{2}\sigma^{2}X^{2q}(s)\frac{d^{2}u(X(s))}{dx^{2}}\right)ds$
$+\int_{0}^{t}\sigma X^{q}(s)\frac{du(X(s))}{dx}dW(s)$
$=u(X_{0})+\int_{0}^{t}Lu(X(s))ds+\int_{0}^{t}\sigma X^{q}(s)\frac{du(X(s))}{dx}dW(s)$ (63)
By the uniform ellipticity of $L$ and the boundedness of $\mathcal{O}$, $\mathbb{P}(\tau_{\mathcal{O}}=\infty)=0$; that is, $\tau_{\mathcal{O}}<\infty$ $\mathbb{P}$-a.s. Let us define the approximated stopping time
$S_{n}=\inf\left\{t\geq 0:\text{dist}(X(t),\,\partial\mathcal{O})<\frac{1}{n}\right\}.$ (64)
Note that with this definition, we have $S_{n}\leq S_{n+1}$ for all $n$, and thus
$S_{n}\leq\tau_{\mathcal{O}}\,\,\,\text{ and }\,\,\,\lim_{n\to\infty}\,S_{n}=\tau_{\mathcal{O}}\,\,\,\mathbb{P}-a.s.$ (65)
Then
$u(X_{n})-u(X_{0})$ $=\int_{0}^{t\wedge S_{n}}Lu(X(s))ds+M_{t}.$ (66)
where $X_{n}=X(t\wedge S_{n})$ and
$M_{t}=\int_{0}^{t\wedge S_{n}}\sigma X^{q}(s)\frac{du(X(s))}{dx}dW(s).$ (67)
On the other hand, by Lemma 2.1, we know that $X^{q}(t)\in\mathcal{M_{T}}$. Thus, from (7) and (32), we can obtain
$\mathbb{E}\left[\sup_{0\leq s\leq t\wedge S_{n}}X^{2q}(t)\right]\leq C\left(1+\mathbb{E}\left[\sup_{0\leq s\leq t\wedge S_{n}}X^{2}(t)\right]\right)<\infty.$ (68)
Then, using the Fubini theorem and the fact that $u\in C^{2}$ in $\mathcal{O}$, we have
$\int_{0}^{t\wedge S_{n}}\mathbb{E}\left[\left(\sigma X^{q}(s)\frac{du(X(s))}{dx}\right)^{2}\right]ds<\infty.$ (69)
Thus, $M_{t}$ is a martingale. Next, applying the product formula to the process $e^{-c(t\wedge S_{n})}\,u(X_{n})$, and then using (66) gives
$d\left(e^{-c(t\wedge S_{n})}\,u(X_{n})\right)$ $=e^{-c(t\wedge S_{n})}\,du(X_{n})+u(X_{n})\,de^{-c(t\wedge S_{n})}+d\langle e^{-c(t\wedge S_{n})},\,u(X_{n})\rangle_{t\wedge S_{n}}$
$=e^{-c(t\wedge S_{n})}\,du(X_{n})-ce^{-c(t\wedge S_{n})}\mathbf{1}_{\{t<S_{n}\}}u(X_{n})dt$
$=e^{-c(t\wedge S_{n})}\mathbf{1}_{\{t<S_{n}\}}\left(Lu(X_{n})-cu(X_{n})\right)dt+e^{-c(t\wedge S_{n})}dM_{t}.$ (70)
Integrating over $0,t\wedge S_{n})]$, and taking the expectation provides
$\mathbb{E}^{x}\left[e^{-c(t\wedge S_{n})}\,u(X_{n})\right]-u(x)$ $=\mathbb{E}^{x}\left[\int_{0}^{t\wedge S_{n}}e^{-cs}\left(Lu(X(s))-cu(X(s))\right)ds\right]$
$=-\mathbb{E}^{x}\left[\int_{0}^{t\wedge S_{n}}e^{-cs}f(X(s))ds\right].$ (71)
Using the fact that
  • 1.
    $\lim_{n\to\infty}\,S_{n}=\tau_{\mathcal{O}}$,
  • 2.
    $\lim_{t\to\infty}\,t\wedge\tau_{\mathcal{O}}=\tau_{\mathcal{O}}$,
and hence, letting first $n\to\infty$ and then $t\to\infty$ in (5), and considering the continuity of the sample paths and $u$, and the boundary condition, the Dominated Convergence theorem gives
$u(x)$ $=\mathbb{E}^{x}\left[\int_{0}^{\tau_{\mathcal{O}}}e^{-cs}\,f(X(s))ds\right],$ (72)
which concludes the proof. ∎